Supplemental Operations Report – April 8th

Supplemental Operations Report 08-Apr-2023

Name of person filing report: Sergii Iakymov

Reason for Report: Routine.

Non-Nominal Systems: Robotic observatory, Hab car, Opportunity
Action taken for non-nominal systems: Observatory is awaiting for new parts to be procured. Hab Car – electrical issues, will be taken to Loa for diagnostic. Opportunity, see below.

Power system:
Solar–SOC Last 24 hours: Max 100% Min 34 % Avg 55.6%. Solar experienced overload on April 6th by excess input power from the Sun. This was noticed by the Crew engineer during work on her project in the ScienceDom. She heard a loud noise and saw a bright flash from one of the battery chargers. The charger went in P05 – Power Limited (Above Voltage Regulation) mode. After consultations with Dr. Rupert, the system was restarted, but it did not change anything. Later during the day Inverters went offline four times in a row without obvious reason and no fault signals. A decision was made to turn on the generator around 7pm to let inverters rest without load. Inverters were restarted next morning and are functioning properly by this moment.
Power System Notes: Generator run time about 11:30pm – 7am. Was not used on April 1st, and from 7pm – 7am on April 6. Hours – 3639.7.

Propane Reading Station Tank: 57%
Propane Reading Director Tank: 78%
Propane Reading Intern Tank: 75%
Propane Reading Generator Tank: 70%

Water (Static Tank) – 389 gallons
Water in GreenHab – 138 gallons
Water (Outpost tank) – 210 gallons
Water in Science Dome: 0 gallons
Hab Toilet Tank emptied: n/a

Sojourner rover used: No
Hours: 190.4
Beginning Charge: 100
Ending Charge: 100
Currently Charging: Yes
Notes on Rovers: All are plugged in, charged, and serviced. Parts needed to be replaced are ordered, see notes below:Opportunity: one battery connector wire corroded out and disconnected, replacement needed.
Curiosity: one battery connector wire is heavily corroded and requires a replacement
Sojourner: two battery connector wires and one positive terminal wire are heavily corroded and require a replacement.

ATV’s Used: (Honda, 350.1, 350.2, 300): None
Reason for use: n/a
Oil Added? No
ATV Fuel Used: 0 Gals
Ethanol Free Gasoline — 0 Gallons
# Hours the ATVs were Used today: 0
Notes on ATV: There was no time for me to get to Richfield (nor good enough weather) to purchase the trailer. I will be doing that remotely and Sergii will pick it up.

HabCar used and why, where? Yes, to the Hanksville landfield and one water run. Battery was not charging from the car and required to be taken out and recharged after one trip to Hanksville. New battery was purchased but it did not fix the issue. Vehicle is scheduled to be taken to Loa for service on April 11th.

Crew Car used and why, where? Yes. By crew from Grand Junction, to Hanksville water runs and supplies, Loa to pick up a new battery for the Hab car..

General notes and comments: nothing to report

Dual Split in Science Dome: off due to warmer weather.

Summary of Internet: Starlink expirinsed 22 min 47 sec outage on April 7th. Current status – nominal.

Summary of suits and radios: Radio #4 was found without a rubber piece of PTT button and was replaced. Headset #5 was malfunctioning and replaced. Suit #8 battery replaced.

EVA COMMS: Nominal

Campus wide inspection, if action taken, what and why: nothing to report

Summary of General Operations: Crew 279 in simulation, all nominal.

Summary of Hab operations: Nominal. First AID kit was resupplied by antibiotic ointment and triangular bandages.

Summary of GreenHab Operations: Two mice were captured and released to the wild. Door is open during sunny days.

Summary of SciDome Operations: Scientific freezer was installed and plugged in.

Summary of any Observatory Issues: Robotic Observatory remains unoperational and awaiting for replacement parts.

Summary of RAM Operations: Nothing to report

Summary of Outpost Operations:

Director’s trailer:
1) Front door lock was replaced due to old key loss;
2) Bathroom faucet was inspected due to lack of pressure in the hot water line. Inspection revealed that it got clogged inside with now way to get it apart. It was replaced. Current status – nominal.
3) Shower head was adjusted. A plastic insert from inside the head was removed, which was preventing a big amount of water from flowing through. This fixed low water flow.
4) Second half of the door for the cat toilet was found and will be installed next week.
5) Trailer was clean prior to my move next week.
Interns trailer:
1) Bathroom faucet has a broken cold water handle and will be replaced next week.
The Trailer parked and is awaiting for the number plate to be installed and photos taken.

Summary of Health and Safety Issues: Section of the tunnel at the ScienceDom was broken by high winds on April 3rd and separated by station personnel.

Questions, Concerns, Supplies needed and Requests: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site